OpenStudy (anonymous):

solve by elimination 6x+2y=-28 6x-2y=-44

OpenStudy (anonymous):

tina... if you add straight down.. what do you get.

OpenStudy (anonymous):

12x 4y -72

OpenStudy (anonymous):

almost. 12x then the y's cancel because 2y - 2y = 0 12x = -72

OpenStudy (anonymous):

ok wasnt sure how that worked

OpenStudy (anonymous):

Now, how do you solve for x?

OpenStudy (anonymous):

12(-6)=-72

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

but then how do i solve for y, the problem is asking for the solution

OpenStudy (anonymous):

good.. so now you have to put that back into one of the equations to find y I will use the first one.

OpenStudy (anonymous):

6x + 2y = -28 6(-6) + 2y = -28 -36 + 2y = -28 2y = 8 y = 4 Answer (-6,4)

OpenStudy (anonymous):

so do u add -36 or subtract it from both sides

OpenStudy (anonymous):

you add 36 because you are getting rid of -36

OpenStudy (anonymous):

the sign before a number ALWAYS goes with the number and you do the opposite.

OpenStudy (anonymous):

-28--36 =8 or -28+-36 makes -64

OpenStudy (anonymous):

-28 + 36 = 8 (-- is +)

OpenStudy (anonymous):

ok thats what i thought so then 2y=8 2(4)=8 y=4

OpenStudy (anonymous):

quick lesson x + 5 = 10 you - 5 because it says + 5 x - 4 = 10 you + 4 because it says - 4 7 + x = 10 you - 7 because you have + 7 -6 + x = 10 you + 6 because you have -6

OpenStudy (anonymous):

2y = 8 because it says 2 times y you divide by 2 2y 8 -- = -- 2 2 the 2's cancel to leave you with y = 4

OpenStudy (anonymous):

(-6,4) is the solution

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you so much